A friend of yours asks you to help him/her draw two bacterial growth curves. He/she gives you the following information: the TA started both cultures (one was labeled culture �X� and the other culture �Y�) at 8:30am using the same overnight culture to inoculate fresh medium. After inoculation of the medium the OD was 0.14 for both cultures and both cultures were incubated at 37 oC. During class the following OD readings were obtained: for culture �X�: 0.19 at 9:50am, 0.28 at 10:10am, 0.29 at 10:30, 0.68 at 10:50 am, and 1.1 at 11:10am. At 11:30am the OD of a 1:10 dilution of the stock culture was 0.29. For culture �Y�: 0.16 at 9:50am, 0.19 at 10:10am, 0.21 at 10:30, 0.24 at 10:50 am, 0.28 at 11:10am and 0.32 at 11:30am. A) Draw the two graphs. B) What was the generation time of these cultures? C) What could be the reason for the different growth pattern? Give a detailed explanation!
Answer #1

A) The growth curves of cultures A and B are given below.

B) Generation time = t/n

Where t = time interval in minutes and n = number of generations.

The time interval taken is 20 minutes for every OD reading

Generation time of cultures A and B =20/7 = 3 minutes

The cultures A and B double in every 3 minutes.

C) Different growth patterns observed in cultures A and B. Culture A has shown typical growth curve with lag pahse, log phase, stationary and decline phases. Culture B is showing constant growth rate since inoculation. This is because, the culture requires even more time for incubation to exhibit exponential growth in log pase, to reach stationary and death or decline phases. The culture B is therefore, in lag pahse and is adjusting itself to the culture conditions.

OD OD 0.35 0.3 0.25 0.2 0.15 0.1 0.05 1.2 0.8 0.6 OD 0.4 0.2 0 8.30am 9.50am 10.10am 10.30am 10.50am 11.10am 11.30am 8.30am 9